OpenStudy (anonymous):

Factor the trinomial, if it's a perfect square. Or state that it's prime if it's prime. \[x ^{2}-13xy+169y ^{2}\]

OpenStudy (anonymous):

Have you heard of the quadratic formula?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

you cannot factor this, nor, i think, can you solve it using the quadratic formula if it was \[x^2-26xy+169y^2\] then it would be a perfect square, because \(136=13^2\) and \(2\times 13=26\)

OpenStudy (anonymous):

You can certainly solve this in the complex plain

OpenStudy (anonymous):

I have an online course and I was thinking this one was not solveable but wasn't sure

OpenStudy (anonymous):

but since you have \(-13xy\) it is not a perfect square i can't land my complex plain

OpenStudy (anonymous):

as nancy reagan used to say "just say no"

OpenStudy (anonymous):

This is the first one I actually had right before putting on here to double check. :) Thank you!!!

OpenStudy (anonymous):

She is doing how it is described in the textbook. She doesn't know why. Its all basic trinomials. Its like basic algebra.

OpenStudy (anonymous):

I don't know how it's done in the textbook because I can't understand it. But yes, it's from an online course so I only have reading explanation and that doesn't quite work. Thank you!

OpenStudy (anonymous):

What about these factors? \[(2x - 13y)(\frac{1}{2}x+13y)\]

OpenStudy (anonymous):

Don't work either.